OpenStudy (anonymous):

Hi can someone help me determine how to solve this ? Consider the points P at which the distance to A (-1,5,3) is twice that of P to B (6,2, -2). Show these points on a sphere and determine its centre and radius.

OpenStudy (anonymous):

@zepdrix Hi, You helped me for a similar problem earlier, Would it be possible to help me with this one please?

OpenStudy (anonymous):

I also have the answer if needed, i just don't know how to get to the solution

OpenStudy (anonymous):

Use the distance formula to get \[ (x+1)^2+(y-5)^2+(z-3)^2=4 \left((x-6)^2+(y-2)^2+(z+2)^2\right) \]

OpenStudy (anonymous):

After expanding and simplifying you get \[ 3 x^2-50 x+3 y^2-6 y+3 z^2+22 z+141=0 \] Which is a sphere

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

So can we find the radius with this ?

OpenStudy (anonymous):

Complete the squares and finish it

OpenStudy (anonymous):

Wait do i find the radius from this equation ? Or does this equation represent coordinates ?

OpenStudy (anonymous):

\[ x^2-\frac{50 x}{3}+\left(\frac{50}{6}\right)^2-\left(\frac{ 50}{6}\right)^2\\+y^2-2 y+4-4+\\z^2+\frac{22 z}{3}+\left(\frac{22}{6}\right)^2-\left(\frac{22}{6 }\right)^2+47=0 \]

OpenStudy (anonymous):

I have to be honest .. I'm not sure what's going on

OpenStudy (anonymous):

You can now read the center and the radius

OpenStudy (anonymous):

The answer my teacehr gave was C(25/3, 1, -11/3) with radius = \[\sqrt{332}/3\]

OpenStudy (anonymous):

So basically we just use the distance formula for the problem ?

OpenStudy (anonymous):

This might be right, I may have made a mistake for the y part. Try to find my mistake

OpenStudy (anonymous):

But why did you write 4 in front of the distance formula ?

OpenStudy (anonymous):

Yes, I did. I should have added 1 and subtracted 1 instead of adding and subtracting 4

OpenStudy (anonymous):

This should have been \[ x^2-\frac{50 x}{3}+\left(\frac{50}{6}\right)^2-\left(\frac{ 50}{6}\right)^2\\+y^2-2 y+1-1+\\z^2+\frac{22 z}{3}+\left(\frac{22}{6}\right)^2-\left(\frac{22}{6 }\right)^2+47=0 \]

OpenStudy (anonymous):

The final answer is \[ \left(x-\frac{25}{3}\right)^2+(y-1)^2+\left(z+\frac{11}{3}\ \right)^2=\left(\frac{\sqrt{332}}{3}\right)^2 \]

OpenStudy (anonymous):

Do you understand it now?

OpenStudy (anonymous):

But why did you write 4 in front of the distance formula ?

OpenStudy (anonymous):

I get the after parts of the problem, It's just the beginning that has me perplexed

OpenStudy (anonymous):

2^2 =4

OpenStudy (anonymous):

We computed the square of the distances. If s = 2t then \[ s^2 = 4 t^2 \]

OpenStudy (anonymous):

Ooohhh I get it now lol

OpenStudy (anonymous):

So the problem is pretty simple

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

I understand it now .. Thanks:D

OpenStudy (anonymous):

YW